OpenStudy (anonymous):

Hello again my friends, I come here for help regarding OCW Scholar Multivariable Calculus, it has left me dumb and confused. If you are keeping up with this course and have covered the topic cross product which is under the Vectors and Matrices section, then you may be of help. I am stuck on the recitation video. I understand everything until the second to last step. How did he get the answer 3 root(2) from looking at the length of the vector <3,0,-3>

OpenStudy (anonymous):

The length of a vector is given by \[(\sqrt{ x^2+y^2+z^2})\] . So 3^2 + 3^2 is 18, which can also be written as \[3\sqrt{2}\] because \[\sqrt{18}=\sqrt{9}*\sqrt{2}=3\sqrt{2}\]

OpenStudy (anonymous):

Correction: Should be 3^2 + 0^2+ (-3)^2 = 18 to follow the formula for vector length entirely